### v1

# Languages Transferred Within the Encoder: On Representation Transfer in Zero-Shot Multilingual Translation

Abstract
Understanding representation transfer in multilingual neural machine translation can reveal the representational issue causing the zero-shot translation deficiency. In this work, we introduce the identity pair, a sentence translated into itself, to address the lack of the base measure in multilingual investigations, as the identity pair represents the optimal state of representation among any language transfers. In our analysis, we demonstrate that the encoder transfers the source language to the representational subspace of the target language instead of the language-agnostic state. Thus, the zero-shot translation deficiency arises because representations are entangled with other languages and are not transferred effectively to the target language. Based on our findings, we propose two methods: 1) low-rank language-specific embedding at the encoder, and 2) language-specific contrastive learning of the representation at the decoder. The experimental results on Europarl-15, TED-19, and OPUS-100 datasets show that our methods substantially enhance the performance of zero-shot translations by improving language transfer capacity, thereby providing practical evidence to support our conclusions.111Partial works done during Zhi Qu’s internship at ASTREC of NICT, Japan. Our codes are available at https://github.com/zhiqu22/ZeroTrans.
Languages Transferred Within the Encoder:
On Representation Transfer in Zero-Shot Multilingual Translation
Zhi Qu† Chenchen Ding†‡ Taro Watanabe† †Nara Institute of Science and Technology, Japan {qu.zhi.pv5, taro}@is.naist.jp ‡National Institute of Information and Communications Technology, Japan chenchen.ding@nict.go.jp
1 Introduction
State-of-the-art neural machine translation systems are adaptable to multilingualism, resulting in a single encoder-decoder model that executes arbitrary translations by adding a tag specified to the target language at the beginning of source sentence Firat et al. (2016); Johnson et al. (2017); Wu et al. (2021). Multilingual neural machine translation (MNMT) is theoretically attractive because zero-shot translations, i.e., translations unseen in training, allow the training of a multilingual model with minimal cost. Unfortunately, the performance of zero-shot translations always lags behind Aharoni et al. (2019); Arivazhagan et al. (2019a); Gu et al. (2019); Yang et al. (2021); Pan et al. (2021); Chen et al. (2023).
Representational analysis in MNMT models guides the improvement of zero-shot translation. However, two different perspectives are demonstrated by prior works: 1) the encoder clusters the different translation representations by the target language Kudugunta et al. (2019); Liu et al. (2021), as shown in Figure 1(a); 2) an ideal encoder is considered to learn only language-agnostic representations, where general cross-lingual features are modeled and transferred among languages Pan et al. (2021); Gu and Feng (2022); Gao et al. (2023), as shown in Figure 1(b). In this work, our investigation proves that these two perspectives are two sides of the same coin, as the encoder aligns translation representations from different source languages within the subspace of the target language.
We introduce the identity pair, which is a pseudo pair translating a sentence to itself, as a proxy for the optimal representation of a language. We then use multiple analytical methods, including singular value canonical correlation analysis (SVCCA) Raghu et al. (2017) and t-distributed stochastic neighbor embedding (t-SNE) van der Maaten and Hinton (2008), to systematically analyze and quantify the representation transfer in MNMT models. We show that transferring representations from the source language to the target language is a major task of the encoder and is entirely accomplished by the encoder. Therefore, the zero-shot translation deficiency arises from the translation representations being entangled with other languages rather than being effectively transferred to the expected target language within the encoder.
Guided by our findings, we propose two methods for the encoder and decoder, respectively, to improve multilingual representations: Low-Rank Language-specific Embedding (LoLE) is applied to bias the representations in the subspaces of target languages at the encoder; and Language-specific Contrastive Learning of Representations (LCLR) is applied at the decoder to isolate representational space across languages. We evaluated the proposed methods on three benchmarks, Europarl-15 Koehn et al. (2005), TED-19 Ye et al. (2018), and OPUS-100 Zhang et al. (2020a); Yang et al. (2021), for two automatic metrics, SacreBLEU Post (2018) and BERTScore Zhang et al. (2020b). The experimental results show that our methods outperform strong baselines in training from scratch because of improved representational transferability. Our methods also perform effectively in fine-tuning, even though pre-trained models are trained by different strategies of language tags, which proves that target language information on the encoder side consistently benefits MNMT.
2 Background
2.1 Multilingual Neural Machine Translation
Johnson et al. (2017); Wu et al. (2021) demonstrated that the training strategy of adding a language tag at the beginning of the input sentence on the encoder side boosts the zero-shot translation capacity of the MNMT model. Given a multilingual corpus that covers a set of languages, a set of their corresponding language tags exists: . For a source-target sentence pair , i.e., and , the training data consists of a pair in form of , where is the language tag of that instructs translation from into language . The model is trained over all pairs in to optimize the following cross-entropy loss:
where is the probability distribution of and represents the model parameters.
2.2 Prior Investigations of MNMT
Two tools are used in prior investigations: 1) The t-distributed stochastic neighbor embedding (t-SNE) van der Maaten and Hinton (2008) converts similarities between a set of vector representations to joint probabilities; 2) The singular value canonical correlation analysis (SVCCA) Raghu et al. (2017) compares the similarity between two vector representations222We follow Liu et al. (2021) to measure SVCCA scores at the sentence level. Appendix A shows how to compute SVCCA scores.. The visualization using t-SNE shows that translation representations, which belong to different source languages but with the same semantics, are aligned at the output of the encoder Pan et al. (2021); Gao et al. (2023), suggesting the ideal encoder is language-agnostic. Specifically, translation representations from different source languages are transferred to a language-agnostic state before being transferred to the target language. On the contrary, the investigation of Kudugunta et al. (2019) based on SVCCA demonstrates that translation representations possess the features of the target language. However, the results derived from measuring SVCCA scores between real translation pairs are disturbed by the lack of base measures. For instance, the translation from English to German, denoted by ende, cannot serve as a base measure for the translation from another language xde, because the representation of ende is not equivalent to the representation of de.
3 Investigating Representation Transfer in MNMT
We conduct preliminary experiments to investigate representations by introducing identity pairs as base measures. Two different datasets, Europarl-15 Koehn et al. (2005); Dabre and Kurohashi (2019) and TED-19 Ye et al. (2018), are applied to verify whether our observations are general. Appendix B introduces the detailed information of those two datasets. Then, following Kudugunta et al. (2019), our investigation is based on Transformer models with 6 encoder and decoder layers, where Appendix C introduces the detailed model settings.
3.1 Identity Pairs
Identity pairs refer to the translation pairs translating from themselves into themselves to present the optimal state of processing language features, i.e., the semantics and syntax of the source sentence by the model. On the encoder side, we derive the representation from a language translating to itself, i.e., , where is the language tag of , with the aim of recovering the source sentence from the hidden representations without inference on the decoder side. We derive the representation in the decoder from the gold translation of .
We use SacreBLEU Papineni et al. (2002); Post (2018) to evaluate the translation quality of 6 identity pairs, which are generated by inference. The scores of enen, dede, and ptpt in Europarl-15 are 73.49, 61.04, and 71.97, which significantly outperform 44.04 of deen, 36.63 of ende, and 46.24 of enpt, respectively. Similarly, enen, trtr, and vivi in TED-19 obtain scores of 72.52, 36.58 and 59.26, which are higher than 34.92 of deen, 14.81 of entr, and 29.78 of envi, respectively. Such high scores in the identity pair are caused by that short sentences are recovered from hidden representations perfectly, and long sentences only have a few changes in word selection. Such evidence suggests that identity pairs can serve as base measures for comparing representations because the identity pair is a proxy for the optimal representation of a language, specifically, xen are expected to be close to enen in the representational space.
3.2 Language Transfer Within the Encoder
Given two languages ① and ②, two sentences of ① and of ② are grouped by three cases when comparing the representations of the encoder: (i) comparing and to show how target language features are encoded333 indicates the identity of ①, i.e., ①①, and indicates a sentence of ② translating to the sentence of ① instructed by the language tag of ①, i.e., ②①.; (ii) comparing and to show how source language features are encoded; (iii) comparing two different identities, and .
The two models trained by Europarl-15 and TED-19 show the same tendency in Figure 2, that is, the language features for ① of (i) consistently increase in both cases involving the central language in Figures 2(a) and 2(c), and non-central languages in Figures 2(b) and 2(d). The target language feature of ① emerges as the primary factor that affects representations at the fifth and sixth layers when the cases of (i), (ii), and (iii) are compared. Therefore, we can conclude that the language features of the representations are transferred to the target side within the encoder. Meanwhile, we also observe that the scores of (iii) are close to or even exceed those of (ii) at some layers both in Figure 2. This proves that the feature of the source language is not the primary factor for representations in the encoder because representations are transferred to the subspaces of target languages. Thus, the comparison between (ii) and (iii) supports the conclusion that language transfer is completed within the encoder.
On the other side, identity pairs also allow the measurement of the alignment of different languages in the target language space through t-SNE. Compared with the sentence-level measurement of Pan et al. (2021); Gao et al. (2023), we measure the alignment of representations at the token level. As shown in Figure 3, semantic similarity causes the representations to cluster together. Moreover, as shown in Appendix H, these representations are not clustered before being processed by the encoder, and the case with different target languages has a higher overall variance. Combined with the finding that the encoder transfers the representation of the source language to the target language, the evidence further suggests that there is no general and cross-lingual state for directly sharing semantic information within the encoder, and the alignment shown in Figure 1(b) occurs in the representational subspace of the target language.
3.3 Entanglements Hindering the Transfer
Although the investigation in Section 3.2 shows that the representations gradually transfer to the target language in a translation pair, the representation spaces of multiple languages may potentially entangle with each other, resulting in the failure of the zero-shot translation Qu and Watanabe (2022). To further illustrate the relationship between different languages, we use t-SNE and BiKDE to visualize the representations at the output of the encoder for the several identity pairs in Europarl-15 and TED-19. Figures 4(a) and 4(d) show that different identity pairs are uniformly distributed in the representational space. This distribution proves again that the encoder is language-specific because each language has an isolated representational subspace.
Compared with identity pairs that represent the ideal capability of the model in processing languages, the distributions plotted in Figures 4(b) and 4(e) reflect the actual capacity for the supervised translation of enx. Figures 4(b) and 4(e) show the distribution of representations in the pairs translating from en, which are similar to that of identity pairs. The difference between identity pairs and supervised language pairs can be attributed to the influence of the source language information, which hinders the full use of the target language information learned by the encoder.
Moreover, the language-specific subspaces cannot be clearly separated for zero-shot translations, as shown in Figures 4(c) and 4(f). Specifically, all representations are entangled around the supervised language pair of xen, which hinders these representations from transferring into the ideal subspaces of the target language. This aligns with Qu and Watanabe (2022) that non-central languages are entangled with the central language, i.e., English, which explains the weakness of zero-shot translation compared with supervised translation, namely, the focal point for improving the transferability of representations is to improve the extent of language transfer in the encoder.
3.4 Language Features in the Decoder
The decoder works on representations with target language features because the source language has already been transferred into the target language subspace (Section 3.2). Therefore, representational entanglements (Section 3.3) also affect the decoder. We further investigate the importance of target language features versus semantics in the decoder. Given two sentences of language ③ and of language ④, the decoder representation of is considered as the base measure. We group two cases: (iv) For each sentence in a test set, we identify the pair with the lowest SVCCA score in the encoder representation to derive a that is distant from . Then, we compare it with the base measure to show the importance of target language features; (v) We compare the base measure and to show the importance of semantics. The two scenarios shown in Figure 5 present the same trend, which is that (iv) maintains high scores despite their semantics being entirely different. At the top layers of the decoder, the gradually increasing difference between (iv) and (v) confirms that the decoder tends to learn the target language specificity Sen et al. (2019). However, Figure 5 shows that, for (iv), a wider interquartile range exists at the bottom layers of the decoder, and its scores are close to those of (v), which implies the weakness of the distinguishing languages in zero-shot translations.
4 Encouraging Representation Transfer
To validate the findings in Section 3, we propose two methods on the encoder and decoder sides, respectively, to improve transferability. Based on the findings in Sections 3.2 and 3.3, improving the extent of language transfer in the encoder can overcome the hindered representations of zero-shot language pairs. We introduce a learnable embedding referred to as Low-rank Language-specific Embedding (LoLE). It serves as biases to force representations to transfer into the target language with negligible cost. Based on the findings in Section 3.4, the capacity for multilingual features is insufficient at the lower layers of the decoder. We introduce Language-specific Contrastive Learning of Representations (LCLR) as an training extra task to regularize the representations to specify the representational boundary for each language.
4.1 Low-Rank Embedding for the Encoder
Let , , be a set of embeddings that correspond one-to-one with the languages in . For a translation , the embedding in corresponding to is denoted by . The hidden representation , where , is extracted before the feed-forward network (FFN) Khandelwal et al. (2021); Xu et al. (2023); Deguchi et al. (2023) at the -th encoder layer. Then, we broadcast to , , and we bias to :
where is the input for the FFN of the -th encoder layer (Figure 6(a)). We execute this biasing at the second-top encoder layer to ensure sufficient capacity for fusing representations and language information, while implicitly allowing lower layers to focus on surface-level information.
The simple language categorization by embedding may lead to a risk of dimensional collapse in the latent space Jing et al. (2022). Thus, we reduce the dimension of to to allow biasing in a low rank, and add it to the head of to simultaneously encourage language transfer and minimize the influence on representations Hu et al. (2021). Figure 6(b) is a spectrum used to illustrate dimensional collapse using a comparison of different in Europarl-15. The spectrum shows that larger dimensions are primarily composed of noise, whereas a dimension that is too small adversely affects the learning of key features.
4.2 Contrastive Learning for the Decoder
Given a training batch, we extract hidden representations from the output of each decoder layer and apply averaged pooling to obtain a fixed-dimensional representation for each sentence. To avoid dimensional collapse Tian (2022); Jing et al. (2022), we also use the head of the representation for contrastive learning, i.e., the vectors in the batch , , .
To prevent a potential invalid training objective in sampling caused by the skewed distribution in a batch, we first define by omitting instances that do not share their target language with any other instance in . For a given instance of , which is the anchor in contrastive learning, we let denote the subset of , including instances with the same target language as , where . Likewise, we define a subset for negative instance . For contrastive learning, we randomly sample the positive instance from and sample negative instances from . Additionally, if , we dynamically clip to . Formally, the objective of LCLR is formulated as
where calculates the similarity of representations using the cosine similarity. The final training objective is the sum of Equations 1 and 2, i.e.,
5 Experiments
5.1 Setup
Datasets
Our experiments comprise three English-centric scenarios, i.e., the training and validation sets only involving translation pairs translating to en or from en. 1) Europarl-15 is a sub-collection of MMCR4NLP Koehn et al. (2005); Dabre and Kurohashi (2019), resulting in 5.3 million training translation pairs. 2) TED-19 collects 19 languages from TED Talks Ye et al. (2018), where the training set comprises 6.5 million pairs in total. 3) OPUS-100 Zhang et al. (2020a); Yang et al. (2021) consists of 95 languages and 109.2 million training pairs. Generally, the validation and test sets of a pair contain 2,000 pairs. More detailed information of those datasets can be found in Appendix B.
| Europarl-15 | TED-19 | OPUS-100 | ||||||||||||||
| BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | |||||||||||
| Method | en | en | zero. | en | en | zero. | en | en | zero. | en | en | zero. | en | en | zero. | zero. |
| Vanilla | 37.49 | 43.39 | 24.65 | 88.50 | 95.71 | 84.27 | 24.53 | 29.67 | 11.98 | 83.77 | 93.54 | 77.74 | 23.37 | 28.30 | 5.04 | 69.98 |
| DisPos | 37.15 | 43.37 | 25.89 | 88.39 | 95.72 | 84.69 | 24.08 | 29.43 | 12.80 | 83.62 | 93.49 | 78.36 | 22.72 | 28.24 | 5.58 | 70.74 |
| TLP | 37.41 | 43.28 | 24.96 | 88.47 | 95.71 | 84.40 | 24.44 | 29.62 | 12.74 | 83.73 | 93.53 | 78.24 | 23.41 | 28.30 | 4.60 | 69.40 |
| SemAli | 37.27 | 43.06 | 25.25 | 88.42 | 95.69 | 84.43 | 23.55 | 28.67 | 13.45 | 83.43 | 93.36 | 78.91 | 22.35 | 28.29 | 6.42 | 72.00 |
| LoLE | 37.62 | 43.50 | 26.09 | 88.51 | 95.72 | 84.81 | 24.39 | 29.72 | 13.20 | 83.74 | 93.54 | 78.65 | 23.15 | 28.28 | 7.92† | 73.32† |
| LCLR | 37.44 | 43.43 | 25.71 | 88.46 | 95.72 | 84.66 | 24.46 | 29.66 | 12.12 | 83.76 | 93.54 | 77.87 | 23.34 | 28.37 | 5.11 | 70.04 |
| BOTH | 37.67† | 43.51 | 26.20† | 88.50 | 95.72 | 84.85† | 24.49 | 29.79 | 13.31 | 83.76 | 93.56 | 78.76 | 23.40 | 28.27 | 7.97† | 73.10† |
Models
In experiments of training from scratch, we implement a Transformer model with 6 encoder and decoder layers. Then, three pre-trained models444Notably, those models are trained by adding a source language tag at the encoder and a target language tag at the decoder. In fine-tuning, we keep their strategy, because it is not conflict with adding extra target information by LoLE. are utilized in our experiments of fine-tuning, including M2M-418M, M2M-1.2B Fan et al. (2020) and mBART50 Liu et al. (2020). Based on the ablation study conducted in the validation set in Europarl-15 and TED-19 shown in Appendix F, we apply LoLE in the fifth encoder layer, set to 128, and set and of LCLR to 64 and 30, respectively. We enlarge to 256 and to 128 for models trained on OPUS-100 and three pre-trained models because they involve more languages. When we solely apply LCLR, we set the position to the bottom decoder layer based on the findings in Section 3.4. When we integrate both LoLE and LCLR into a model, we relocate LCLR to the second-bottom decoder layer because of the improved language features of the encoder representations. More detailed settings of models and training details can be found in Appendix C.
Baselines and Evaluation
The vanilla Transformer Vaswani et al. (2017); Johnson et al. (2017) is a baseline in experiments, denoted by Vanilla. Additionally, three representative methods are reproduced in experiments of training from scratch with our datasets and settings: 1) SemAli Pan et al. (2021), which is a state-of-the-art MNMT model without fine-tuning that uses contrastive learning in the encoder output for semantic alignment; 2) DisPos Liu et al. (2021), which removes one residual connection in the encoder to weaken the source language information; and 3) TLP Yang et al. (2021), which has an extra loss introduced to identify the language of the decoder’s output representation. In the evaluation, we set the beam size to 4 in inference. We not only use SacreBLEU Papineni et al. (2002); Post (2018) to evaluate the quality of inferences but also report BERTScore Zhang et al. (2020b) of inferences at the representation level. We also conduct statistical significance testing Koehn (2004) using paired bootstrap resampling with 1,000 iterations and 0.5 resampling ratio, consequently, the case of 0.05 means that the difference is significant.
| BLEU | BERTScore | |||||
| Method | en | en | zero.* | en | en | zero* |
| M2M-418M | 21.88 | 26.43 | 14.51 | 82.52 | 93.25 | 79.26 |
| F.T. | 26.68 | 32.95 | 17.46 | 84.47 | 94.30 | 80.79 |
| LoLE | 26.81 | 33.16 | 17.52 | 84.51 | 94.31 | 80.84 |
| LCLR | 26.81 | 33.67† | 17.65 | 84.47 | 94.40 | 80.88 |
| BOTH | 26.83† | 33.63† | 17.68† | 84.49 | 94.38 | 80.90† |
| M2M-1.2B | 24.32 | 28.94 | 15.95 | 83.17 | 93.72 | 79.75 |
| F.T. | 27.71 | 34.97 | 18.48 | 84.71 | 94.53 | 81.14 |
| LoLE | 28.29† | 34.12 | 18.67† | 84.91† | 94.48 | 81.26† |
| LCLR | 28.26† | 34.54 | 18.64 | 84.90† | 94.50 | 81.22 |
| BOTH | 28.37† | 34.59 | 18.69† | 84.92† | 94.51 | 81.23 |
| mBART50 | 25.28 | 33.50 | 6.92 | 83.93 | 94.43 | 72.91 |
| F.T. | 27.17 | 33.96 | 5.58 | 84.64 | 94.36 | 72.96 |
| LoLE | 27.19 | 33.93 | 7.28 | 84.60 | 94.37 | 73.86 |
| LCLR | 27.07 | 34.02 | 9.69† | 84.59 | 94.38 | 75.31† |
| BOTH | 27.36 | 34.04 | 9.55† | 84.66 | 94.36 | 75.55† |
5.2 Results
Table 1 shows the experimental results of training from scratch. Our proposed methods always perform best in comparison with strong baselines, or are close to the best method. In supervised translations of three benchmarks of Europarl-15/TED-19/OPUS-100, LoLE shows divergent results of 0.13/-0.04/-0.22 on enx and 0.11/0.05/-0.02 on xen. Similarly, LCLR shows diverse results of -0.05/-0.01/-0.03 and 0.04/-0.01/0.07, respectively. Then, Both achieves the results of 0.18/-0.02/0.03 on enx and 0.12/0.12/-0.03 on xen. The results suggest that our model is comparable to Vanilla on supervised translations. Therefore, given that the improvement of zero-shot translation potentially reduces the quality of supervised translations Gu et al. (2019); Zhang et al. (2020a); Liu et al. (2021), our methods are stronger in avoiding the degeneration of supervised language pairs compared with other methods. In zero-shot translations, BOTH outperforms Vanilla 1.55/1.33/2.93 for BLEU and 0.58/1.02/3.12 for BERTScore. In an assessment of the qualities individually, LoLE achieves gains of 1.53/1.22/2.85 for BLEU and 0.54/0.91/3.34 for BERTScore, and LCLR improves 1.06/0.14/0.09 scores for BLEU and 0.39/0.13/0.06 scores for BERTScore. The results indicate that our proposed methods are orthogonal. Moreover, the significant testing shows that our methods substantially improve the zero-shot translation capacity in training from scratch.
Table 2 shows the experimental results of fine-tuning. For M2M-418M, compared with F.T., our methods obtain up to 0.15/0.72/0.22 for BLEU scores and 0.04/0.10/0.11 for BERTScore in enx, xen and zero-shot translations, respectively; For M2M-1.2B, the gain is up to 0.66/-0.38/0.21 for BLEU scores and 0.21/-0.02/0.12 for BERTScore; For mBART50, the gain is up to 0.19/0.08/4.11 for BLEU scores and 0.02/0.02/1.69 for BERTScore. Those scores show that the improvement on M2M is marginal compared with training from scratch. This derives M2M is trained by interconnected translation pairs instead of an English-centric dataset, which results in the robust transferability of multilingual representations. However, the degeneration on F.T. of mBART50 shows that fine-tuning drastically influences the zero-shot translation capacity. For instance, the BLUE scores of frvi decrease to 11.84 from 20.57 and frzh increase to 13.52 from 1.90, but our model obtains 18.47 and 17.19, respectively. Such results and the significant testing indicate again the advantage of our proposed methods in improving multilingual representations for zero-shot translation capacity.
6 Discussion
6.1 Correlation between Representational Disentanglements and Improvements
Table 1 shows the overall results by taking averages across all language pairs, which may overlook pair-specific tendencies. Therefore, we group Europarl-15 by the language families and report the average scores of translating from one language family to another. Figure 7(a) shows the difference in BLEU scores between our models and Vanilla. As shown in Figure 7(b), we also compute the SVCCA scores between the identity of the non-central language and the identity of the central language at the encoder’s output and group them in the same manner. Given the similar distribution in Figure 7, we conduct Pearson correlation analysis Pearson (1896) of all language pairs instead of language families in Europarl-15, and we compute the coefficients and -values of Pearson correlation by target languages to maintain fairness. We observe two key points: 1) The coefficient and -value of en are -0.087 and 0.76, respectively. This result suggests that there is no statistical correlation, which is predictable because xen is not affected by representational entanglements. 2) The coefficient and -value of non-central languages are in the ranges of 0.585 to 0.855 and 4e-5 to 0.021, respectively. In more detail, the mean values are 0.770 and 0.002 and the variances are 0.04 and 3e-5, respectively. This analysis proves that the degree of representational disentanglement positively correlates with the improvement of zero-shot translations.
6.2 Analysis of Improved Representation
To further verify our hypothesis, we measure representation transfer in the model incorporating our proposed methods. As shown in Figures 8(a) and 8(b), both scenarios exhibit improvements on (i). Meanwhile, Figures 8(c) and 8(d) indicate that the entanglement of representations among languages at the encoder is resolved. The evidence suggests that LoLE effectively enhances representation transfer in the encoder. Additionally, (ii) and (iii) in Figures 8(a) and 8(b) also achieve higher scores at lower layers of the encoder, which suggests that LoLE indeed makes lower layers of the encoder focus on surface-level information. By contrast, as shown in Figures 8(e) and 8(f), the more stable trend of (iv) in both scenarios suggests that LCLR can improve the capacity of lower layers of the decoder to distinguish languages to improve zero-shot translations. In addition, Appendix G provides the representational analysis for fine-tuning models, which proves that target language features are consistently beneficial in the encoder.
7 Related Works
Although representational analysis can guide improving the zero-shot translation capacity in MNMT, prior analyses have shown seemingly contradictory conclusions. Kudugunta et al. (2019) found that the encoder clusters representations of different translations based on target language features, suggesting that the encoder is language-specific. However, this conclusion is indirect because comparing real translation pairs would accumulate errors, even though Gu et al. (2019); Liu et al. (2021); Qu and Watanabe (2022) can support the validity. Our work extends this conclusion by using identity pairs to show that representations of source languages are transferred to the target subspace. Notably, Zhang et al. (2021); Pires et al. (2023) reached similar conclusions by incorporating learnable weights in additional language-specific modules. Another viewpoint is that an ideal encoder is considered to learn only language-agnostic representations Pham et al. (2019); Zhu et al. (2020); Pan et al. (2021); Gu and Feng (2022); Gao et al. (2023). Our representational analysis demonstrates that the general cross-lingual state does not exist in the encoder and the alignment occurs within the representational subspace of the target language. Additionally, our analysis supports that delimiting the language-specific representation boundaries at the decoder side is also important Yang et al. (2021); Chen et al. (2023).
8 Conclusion
We systematically investigated the representational issue of zero-shot translation deficiency in multilingual neural machine translation models. Our analyses show that the encoder transfers translation representations from the source language to the target language, and aligns different languages at the target language subspace. We applied engineering practices to verify our findings by proposing two orthogonal methods, which substantially improve the zero-shot translation capacity. Thus, our findings are significant for guiding the improvement of the transferability of multilingual representations.
9 Limitations
This work has two limitations. First, the identity pair is a proxy of language-specific representations, which are based on bi-directional training, i.e., each non-central language appears in the encoder and decoder together. Therefore, we designed an additional study to investigate the impact by retraining a model by eliminating nlen and enit so that it and nl appear only in the encoder and decoder, respectively based on the analysis in Section 3.2. Then, we conducted a comparison by taking de as the middle language to perform the role of identity pairs in analysis. As shown in Figure 9, the target language features keep the same trend as shown in Section 3.2 to support our conclusion again, but the influence of source language features increases relatively.
The second limitation is our investigation is based on adding a language tag specified to the target language at the beginning of the source sentence for the encoder. Although this is the de facto MNMT training strategy Johnson et al. (2017); Aharoni et al. (2019); Arivazhagan et al. (2019a); Gu et al. (2019); Pham et al. (2019); Wu et al. (2021); Yang et al. (2021); Pan et al. (2021); Qu and Watanabe (2022); Chen et al. (2023); Gu and Feng (2022); Gao et al. (2023), the current pre-trained models Fan et al. (2020); Liu et al. (2020); Team et al. (2022) are not based on the strategy. Our methodologies are also effective in mBART50 Liu et al. (2020) and M2M Fan et al. (2020) because they are based on adding a source language tag for the encoder, where we can measure the effectiveness of improving the target language information. However, NLLB Team et al. (2022) is based on two target language tags at the encoder and the decoder, respectively, whose representation transfer is not clear from the conclusions of this work. Thus, our future work is to investigate the representation transfer of NLLB.
Appendix A Sentence-level SVCCA Score
We use SVCCA Raghu et al. (2017) to measure representation similarity in MNMT Kudugunta et al. (2019). We follow the approach of Liu et al. (2021) so that similarity is measured at the sentence level to ensure that each score is computed on equivalent features without the influence of other sentences in the set.
Based on the definition of Section 2.1, we denote hidden representations of a sentence by }, where , equals to the length or from either the encoder or decoder, and is the model dimension. Additionally, the practical length is when is fed into the encoder because the encoder receives the input concatenated by and . Then, we derive the sentence-level representation using average pooling , which mainly represents the language features and semantics of the source sentence rather than syntactic information because positional information is reduced.
Given and derived from two sentences, SVCCA first performs singular value decomposition on their averaged representations to obtain subspace representations and , where noise is reduced Saphra and Lopez (2019). Then we perform canonical correlation analysis Hardoon et al. (2004) to determine and . Formally, we compute correlation between and as
where indicates the inner product. We use to represent the similarity of two sentences. Finally, we compute the set-level score by taking the average scores of all sentences over the set.
Appendix B Detailed Information of Datasets
This work involves three datasets, i.e., Europarl-15, TED-19, and OPUS-100, where Europarl-15 and TED-19 are used in preliminary experiments.
Europarl-15 is collected from MMCR4NLP, which has high-quality sentences and each sentence in a language is one-to-one corresponding to other languages, i.e., all language-specific sets have parallel semantics Koehn et al. (2005); Dabre and Kurohashi (2019), including 15 European languages from 4 language families. Specifically, Germanic includes en, de, nl, da, Romance includes es, pt, it, ro, Slavic includes sl, bg, pl, cs, and Uralic includes fi, et, hu. The training and validation sets cover 28 supervised language pairs where English is the central language used to bridge the non-central languages. The test set consists of all language pairs, including 182 zero-shot language pairs in addition to supervised language pairs. Each pair in the training set comprises 189,310 sentences. The validation and test sets for each pair contain 2,000 sentences, and we measure SVCCA scores in the test sets.
In contrast to the aligned corpus, TED-19 consists of 19 languages, including en, ar, he, ru, ko, it, ja, zh, es, nl, vi, tr, fr, pl, ro, fa, hr, cs, de, which belong to various language families without parallel semantics, from TED Talks Ye et al. (2018). Each pair contains 103,093 to 214,111 sentences in training. The training set comprises 6,551,456 sentences in total, and the validation and test sets for each pair contain 2,000 sentences. Because of the unparallel semantics of TED-19, we align ar, he, zh, hr, vi, ja to obtain 967 sentences for measuring SVCCA scores. In addition, the reason why the number of languages is 19 is that, first, TED Talks have 20 high-resource languages, which are supported in M2M Fan et al. (2020) and mBART50 Liu et al. (2020). However, the tokenization of th is problematic, resulting in deprecating th.
OPUS-100 consists of 95 languages, 188 pairs, and 109.2 million samples in total Zhang et al. (2020a); Yang et al. (2021), where 90 pairs comprise 1 million samples and 56 pairs have more than 0.1 million samples. Different from Yang et al. (2021), we do not use the provided valid set for zero-shot translation pairs in training to avoid biases when assessing the transferability of multilingual representations.
Appendix C Detailed Settings of Models
We use the Transformer Vaswani et al. (2017) from Fairseq Ott et al. (2019) as the model. For the configuration of models trained on Europarl-15 and TED-19, we follow Kudugunta et al. (2019) to set 6 encoder and decoder layers. In addition, we investigate models in Appendix E with 8 and 10 encoder layers. We adopt a shared vocabulary trained by SentencePiece Kudo and Richardson (2018) with 50,000 tokens for both the encoder and decoder. The model consists of 4 attention heads, embedding size of 512, inner size of 1024, dropout rate of 0.2, maximum learning rate of 0.0005 with the inverse square root schedule and 4,000 warmup steps, and label smoothing rate of 0.1. We set the batch size to 8,000 tokens per GPU, apply Adam Kingma and Ba (2017) as the optimizer, and set temperature sampling with Arivazhagan et al. (2019b). We train the model with 60 epochs for Europarl-15 and 30 epochs for TED-15, and finally average the top 5 checkpoints using the loss on the validation set. Compared with the basic configuration, the models trained on OPUS-100 have 8 attention heads, embedding size of 512, inner size of 2048, dropout rate of 0.1, and shared vocabulary size of 64,000. We train the model of OPUS-100 for 400,000 update steps with a batch size of 8,000 tokens per GPU for OPUS-100 and directly use the best checkpoint selected using the loss on the validation set. Furthermore, models with Europarl-15 and TED-19 are trained on 8 NVIDIA V100 GPUs, and models with OPUS-100 are trained on 4 NVIDIA A6000 GPUs by setting –update-freq to 2 in Fairseq to simulate 8 GPUs.
Three pre-trained models are utilized in our experiments. The first is M2M-418M Fan et al. (2020), a model trained on standard multilingual translation tasks, that supports translation across 100 languages. It is based on Transformer architecture, configured with 12 encoder and decoder layers, embedding size of 1024, inner size of 4096, and vocabulary size of 128,112, which results in a total of 418 million parameters. The second model, M2M-1.2B Fan et al. (2020), extends the M2M-418M by enlarging the number of layers to 24 and the inner size to 8192, which culminates in 1.2 billion parameters. The last model is mBART50 Liu et al. (2020), which was trained on monolingual corpora across 50 languages following Lewis et al. (2019) and preliminarily fine-tuned for multilingual translation tasks. It shares the same parameter setup as M2M-418M with a vocabulary size of 250,053, which consists of 611 million parameters. We conduct experiments on TED-19 because all covered languages are supported by these models.
Appendix D Detailed Introductions of Figure 1
In fact, Figure 1(a) corresponds to the last subfigure of Figure 10 to show the linguistic affinity between translations from English to other languages, denoted by enx. Specifically, Figure 10 shows the layer-wise states of the encoder, and Figure 1(a) (Figure 10(f)) demonstrates the state at the output of the encoder. We employ sklearn.manifold.SpectralEmbedding, referring to https://scikit-learn.org, to visualize the similarities computed by SVCCA (Appendix A) for every layer in the encoder. Then, we can find that representations at all encoder layers have certain clusters influenced by the families of the target languages, and the clusters become more distinct as the depth of the encoder layers increases. This suggests that the transfer of representations to the target language begins as early as the first layer of the encoder, with gradual strengthening through further layers. Meanwhile, this finding, i.e., even the initial encoder layers capture target language features, complements prior works Kudugunta et al. (2019); Pires et al. (2023).
On the other hand, we follow Pan et al. (2021) and Gao et al. (2023) to measure the alignment of encoder representations between the identity of en and source languages from different families to English using t-distributed stochastic neighbor embedding (t-SNE) van der Maaten and Hinton (2008) and bivariate kernel density estimation (BiKDE) Wand and Jones (1993). As shown in Figure 11, representations from the four language families are all highly aligned with the identity pair of enen, where the common feature of those translations is the parallel semantics. Thus, this proves that the encoder semantically aligns different translations. However, the deep discussion should be referred to Section 3.2.
Appendix E Investigation on Models with 8 and 10 Encoder Layers
Figure 12 shows the scenarios of 8 and 10 encoder layers to confirm our conclusion, where all the cases for (i) consistently reflect the same tendency. Together with the increased number of layers, the increased final scores of (i) indicate the stronger target language features of ① presented for representations. Meanwhile, we also observe that the scores of (iii) are close to or even exceed those of (ii) at some layers both in Figures 2 and 12. This proves that the feature of the source language is not the primary factor for representations in the encoder because representations are transferred to the subspaces of target languages. Thus, this further investigation supports our conclusion that language transferred within the encoder is general.
Appendix F Selecting Hyper-Parameters
We conduct ablation studies on the validation set of Europarl-15 to select hyper-parameters for LoLE and LCLR, which are used in Section 5.1. Figure 13(a) shows that LoLE performs optimally with the dimension of 128, which corroborates our hypothesis in Section 4.1. Figure 13(b) indicates that LoLE performs the best at the fifth layer and degrades significantly at lower layers, which aligns with our assertion in Section 3.2 that lower layers of the encoder are more correlated with the source language, and enhances the language transfer benefits of transferability (Section 4.1). Figure 13(c) is consistent with the theory of contrastive learning in which full dimensions lead to collapse Jing et al. (2022). Figure 13(d) demonstrates that, as the position constructed by LCLR increases, the scores decrease, which lends support to our analysis in Sections 3.4 that the instability of decoder representations primarily manifests at lower layers, which also explains the weakness of TLP because improving the capacity of distinguishing languages is redundant for the decoder’s top layer. We also conduct an ablation study for hyperparameter for LCLR with a dimension of 64 at the bottom decoder layer. The results are shown in Figure 13(e), with an empirically optimal .
Appendix G Analysis of Improved Representation for Fine-tuning Pre-trained Models
| Pairs | Model | Method | (i) | (ii) | (iii) | |
| Encoder Side | ① of zh ② of ar | M2M | F.T. | 79.66 | 100.0 | 79.66 |
| LoLE | 79.52 | 98.75 | 77.76 | |||
[…]
| LoLE | 63.52 | 97.90 | 61.66 | |||
[…]
| LoLE | 80.54 | 98.27 | 79.88 | |||
[…]
| LoLE | 70.46 | 97.61 | 67.04 | |||
| Decoder Side | ① of ja ② of he | M2M | F.T. | 99.80 | 92.01 | 92.65 |
| LoLE | 99.73 | 89.81 | 90.66 | |||
[…]
| LoLE | 98.64 | 90.07 | 88.49 |
Section 6.2 is the representational analysis for models, which are trained from scratch with proposed LOLE and LCLR. We also show the representational analysis for fine-tuning pre-trained models.
Given the positive correlation shown in Section 6.1, we compute SVCCA scores in the same way as done in Section 3.2 and show the results in Table 3. Unlike Section 3.2, we equally consider the encoder and decoder because the encoder is only related to the source language and does not transfer representations to the target language in the training strategy of M2M Fan et al. (2020) and mBART50 Liu et al. (2020). Additionally, the different training strategy is the primary reason that F.T. shows the same scores in (i) and (iii) and keeps 100.0 in (ii). Alternatively, although the scores of (i), which reflect target language features, decrease in our methods, the scores of (ii) and (iii) also decrease. As a result, the differences between the scores of (i), (ii), and (iii) increase, that is, the relative importance of target language features increases. This result proves our statements in Sections 3.3 and 6.1 again that target language features are consistently beneficial in the encoder. On the other hand, the decoder side shows the same tendency as the encoder side. This fits our motivation in Section 4.2 to further improve the discriminating ability of lower layers of the decoder, although the training strategy of M2M and mBART50 has already provided a high capacity in discrimination for the decoder.
Appendix H Token-level Alignments in Other Cases
First of all, the English sentence for semantic analysis in Figures 3 and 14 is: By the end of this year, there will be nearly a billion people on this planet that actively use social networking sites. Compared with the discussion in Section 3.2, token-level representations are not aligned at the embedding layer, and are relatively aligned in the case of using the identity pairs, where the degree of divergence is substantially higher than the case of Figure 3.